A wheel of 4cm rad rotates on stationary axis it uniformly speed from rest in a speed of 900 rev/m in a time of 10secs. Find (a) the constant angular acceleration of the wheel (b) tangential acceleration of a point on its rim
1
Expert's answer
2021-07-20T09:45:54-0400

Given:

"n_0=0\:\rm rev/s"

"n=900\:\rm rev/s"

"r=0.04\:\rm m"

"t=10\:\rm s"

(a)

"\beta=\frac{\omega-\omega_0}{t}=2\pi\frac{n-n_0}{t}"

"=2\pi*\frac{900-0}{10}=565\:\rm rad/s^2"

(b)

"a_{\tau}=\beta r=565*0.04=22.6\:\rm m/s^2"